The following income statement was drawn from the records of Campbell Company, a merchandising firm: CAMPBELL COMPANY Income Statement For the Year Ended December 31, 2018 Sales revenue (5,500 units x $163) Cost of goods sold (5,500 units x $83) Gross margin Sales commissions (10% of sales) Administrative salaries expense Advertising expense Depreciation expense Shipping and handling expenses (5,500 units x $2) Net income $ 896,500 456,500 440,000 (89,650) (83,000) (33,000) (49,000) (11,000) $ 174,350 Required a. Reconstruct the income statement using the contribution margin format. b. Calculate the magnitude of operating leverage c. Use the measure of operating leverage to determine the amount of net income Campbell will earn if sales increase by 20 percent.Required A Req B and C Reconstruct the income statement using the contribution margin format. CAMPBELL COMPANY Income Statement For the Year Ended December 31, 2018 Less: Variable costs Less: Fixed costsComplete this question by entering your answers in the tabs below. Required AReq B and C Calculate the magnitude of operating leverage. Use the measure of operating leverage to determine the amount of net income Campbell will earn if sales increase by 20 percent. (Round intermediate and Operating leverage answer to 2 decimal places.) b. Operating leverage C. Net income times Required A Req B and C
Answer #1
A
Sales revenue 896500
Less: Variable costs
Cost of goods sold 456500
Sales Commissions 89650
Shipping and handling expenses 11000
Contribution margin 339350
Less: Fixed costs
Administrative salaries expense 83000
Advertising expense 33000
Depreciation expense 49000
Net income 174350
B
Operating leverage = Contribution margin/Net income
Operating leverage = 339350/174350= 1.95 times
C
Net income = 174350+(174350*1.95*20%)= $242347